Two objects (X and Y) are placed at a distance of 0.200 m from each other. The charge on X is 1.38 x 10-6 C and the  charge on Y is 2.48 x 10-6 C. What is the value of the electrical force (in Newtons) between the two objects? 



1
Expert's answer
2021-01-04T14:40:21-0500

"F=k\frac{q_1q_2}{r^2}=9\cdot10^9\cdot\frac{1.38\cdot10^{-6}\cdot2.48\cdot10^{-6}}{0.2^2}=0.77(N)"